# Maximal discs of Weil-Petersson class in $\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}$

### v1

@@ 4.2. Symplectic forms
For a harmonic diffeomorphism $F_{\pm}:\Sigma\to\mathbb{D}_{\pm}:=(\mathbb{D},z_{\pm})$ in the diagram (4.14), we denote by $(\mu_{+},\mu_{-})$ the Beltrami differential representing $(\mathbb{D},z_{\pm})$ respectively. As explained in the subsection 2.2, let us fix an orthonormal basis of in $H^{-1,1}(\mathbb{D})$ and we obtain the corresponding orthonormal basis for $T_{[\mu]}T_{0}(1)$ by the right translations, that is, given by $D_{0}R_{[\mu]}(H^{-1,1}(\mathbb{D}))$ for any $[\mu]\in T_{0}(1)$. We denote by $\{\nu_{\pm,k}\}_{k\in\mathbb{Z}}$ the orthonormal basis of the tangent space $T_{[\mu_{\pm}]}T_{0}(1)$ obtained in this way respectively. For a harmonic Beltrami differential $\nu_{\pm,k}$ on $(\mathbb{D},z_{\pm})$, we solve the Beltrami differential equation
(4.19) $(h^{\epsilon}_{\pm})_{\bar{z}}=\epsilon\nu_{\pm,k}(h^{\epsilon}_{\pm})_{z}\qquad\text{where}\quad z=z_{\pm}$
to obtain a quasi-conformal map $h^{\epsilon}_{\pm}:\mathbb{D}_{\pm}\to\mathbb{D}_{\pm}^{\epsilon}$. Note that $h^{\epsilon}_{\pm}$ and $\mathbb{D}^{\epsilon}_{\pm}$ depends on $k$, but we do not denote it for simplicity of notations. Now, for given two above discs $\mathbb{D}^{\epsilon}_{\pm}$, as the way of the inverse of the map $\mathrm{Mess}$, we construct a maximal disc $\Sigma^{\epsilon}$ in $\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}$ such that there exist harmonic maps $F_{\pm}^{\epsilon}:\Sigma^{\epsilon}\to\mathbb{D}^{\epsilon}_{\pm}$. Then, there exists a unique quasi-conformal map $f^{\epsilon}:\Sigma\to\Sigma^{\epsilon}$ such that
@@
This can be easily follows from the fact that the maximal disc $\Sigma^{\epsilon}$ is the graph of the map $F^{\epsilon}_{+}\circ(F^{\epsilon}_{-})^{-1}$, and that a point in $\Sigma^{\epsilon}$ is determined by its two coordinates given by $F^{\epsilon}_{\pm}$. Note that the map $f^{\epsilon}$ and the maximal disc $\Sigma^{\epsilon}$ depend on $k$. Now we put $\nu_{k}$ to be the Beltrami differential on $\Sigma$ such that $(f^{\epsilon}_{k})_{\bar{z}}=\epsilon\nu_{k}(f^{\epsilon}_{k})_{z}$. Here $f^{\epsilon}_{k}$ denotes the quasi-conformal map $f^{\epsilon}$ corresponding $k$.
Now we define the variation of $\mu$ along $\nu_{k}$ by
@@
where the second equality follows from (4.16). By (4.6), we can see that $F^{*}_{\pm}(\nu_{\pm,k})$ is determined by $\nu_{k}$ and $\mu_{F}=\mu_{F_{+}}$. We also define the variation of $\Phi$ along $\nu_{k}$ by
@@
where the second equality follows from (4.15).
@@
The differential of $\mathrm{Mess}:T^{*}T_{0}(1)\to T_{0}(1)\times T_{0}(1)$ is an isomorphism.
@@
By the proof of Theorem 3.10, the map $\mathrm{Mess}$ is one-to-one and onto, hence its inverse $\mathrm{Mess}^{-1}$ is well-defined. Now, we want to show that the differential of $\mathrm{Mess}^{-1}$ is an isomorphism. First, we observe that the differential of $\mathrm{Mess}^{-1}$ is given by (4.21) and (4.22). In order to analyze these maps, we decompose a tangent space $T(T_{0}(1)\times T_{0}(1))$ at $z_{\pm}|_{\mathbb{S}^{1}}$ into the diagonal tangent spaces and the anti-diagonal tangent space. Then, along the diagonal tangent space, we have $F^{*}_{+}(\nu_{+,k})=F^{*}_{-}(\nu_{-,k})$. Hence, the component of the differential of $\mathrm{Mess}^{-1}$ given in (4.22) vanishes over this subspace. If the component of the differential of $\mathrm{Mess}^{-1}$ given in (4.21) has a nontrivial kernel, then we should have
$F^{*}_{+}(\nu_{+})+F^{*}_{+}(\bar{\nu}_{+})\mu_{F}^{2}=0$
for a harmonic Beltrami differential $\nu_{+}$. Combining this with its conjugated equality, we have $F^{*}_{+}(\nu_{+})(1-|\mu_{F}|^{4})=0$, which can not happen since $|\mu_{F}|<1$ and $F_{+}$ is a diffeomorphism. Hence, the component of the differential of $\mathrm{Mess}^{-1}$ given in (4.21) is injective. To prove that this component is onto over the diagonal tangent space, pick a harmonic Beltrami differential $\nu$ representing a tangent vector $\delta\mu$, then there exists a Beltrami differential $\nu_{+}$ such that
$\nu-\bar{\nu}\mu_{F}^{2}=(1-|\mu_{F}|^{2})F_{+}^{*}(\nu_{+})$
by the fact $F$ is a diffeomorphism. This implies that
@@
Since $\nu$ can be written in terms of $\nu_{k}$’s, the same is true for $\nu_{+}$ in terms of $\nu_{+,k}$’s. Hence, the component is onto over the diagonal tangent space. In a similar way, we can show that the differential of $\mathrm{Mess}^{-1}$ given in (4.22) is bijective over anti-diagonal tangent space. Hence, these together means that the differential of $\mathrm{Mess}:T^{*}T_{0}(1)\to T_{0}(1)\times T_{0}(1)$ is an isomorphism. This completes the proof. ∎
By Proposition 4.8, we can see that $\{\delta_{k}\mu\}_{k\in\mathbb{Z}}$ and $\{\delta_{k}\Phi\}_{k\in\mathbb{Z}}$ together are basis of the tangent bundle of $T^{*}T_{0}(1)$ at $(\mu,\Phi)$. Using this fact, we define the symplectic form $\omega_{\mathrm{C}}$ on the cotangent bundle $T^{*}T_{0}(1)$ over an open neighborhood $V_{0}$ of the origin in $T_{0}(1)$ by
(4.23) $\omega_{\mathrm{C}}=\int_{\Sigma}\Big{(}\sum_{k\in\mathbb{Z}}\sqrt{-1}\delta_{k}\Phi\wedge\delta_{k}\mu-\sqrt{-1}\delta_{k}\overline{\Phi}\wedge\delta_{k}\bar{\mu}\Big{)}\,d^{2}z.$
Note that the wedge product $\wedge$ is taken for the vectors $F^{*}_{\pm}(\nu_{\pm,k})$ which are vector terms on the right hand side of (4.21) and (4.22). This formula defines the symplectic form on the coordinate chart $V_{0}$. This symplectic form also extends to other charts $V_{\mu}$ by right translation. The definition of $w_{\mathrm{C}}$ involves the choice of basis $\{\nu_{\pm,k}|\,k\in\mathbb{Z}\}$ a priori. But, we will show that $\omega_{\mathrm{C}}$ does not depend on this choice (see Remark 4.11).
@@
The following equality holds
(4.24) $\begin{split}\omega_{\mathrm{C}}=&-\frac{\sqrt{-1}}{2}\int_{\Sigma}(1-|\mu_{F_{+}}|^{2})\,e^{\phi}\\ &\qquad\qquad\cdot\,\sum_{k\in\mathbb{Z}}\Big{(}F^{*}_{+}(\nu_{+,k})\wedge F^{*}_{+}(\bar{\nu}_{+,k})-F^{*}_{-}(\nu_{-,k})\wedge F^{*}_{-}(\bar{\nu}_{-,k})\Big{)}\,d^{2}z.\end{split}$
@@
$\qquad\qquad\qquad\qquad\qquad\qquad\quad\wedge\Big{(}(F^{*}_{+}(\bar{\nu}_{+,k})+F^{*}_{-}(\bar{\nu}_{-,k}))+(F^{*}_{+}({\nu}_{+,k})+F^{*}_{-}({\nu}_{-,k})\bar{\mu}_{F}^{2}\Big{)}$
@@
Recall that the basis $\{\nu_{\pm,k}\}_{k\in\mathbb{Z}}$ is an orthonormal basis at $T^{*}_{[\mu_{\pm}]}T_{0}(1)$ and that the Weil-Petersson symplectic form $\omega_{\mathrm{WP}}$ on $T^{*}_{[\mu_{\pm}]}T_{0}(1)$ has the following expression,
@@
The map $\mathrm{Mess}:T^{*}T_{0}(1)\to T_{0}(1)\times T_{0}(1)$ is a symplectic diffeomorphism such that
@@
Here $\mathrm{Mess}_{\pm}:=\pi_{\pm}\circ\mathrm{Mess}$ where $\pi_{\pm}$ denotes the projection map from $T_{0}(1)\times T_{0}(1)$ onto the first or second factor respectively.
@@
By Proposition 4.9,
(4.27) $\begin{split}\omega_{\mathrm{C}}=&-\frac{\sqrt{-1}}{2}\int_{\Sigma}(1-|\mu_{F}|^{2})\,e^{\phi}\\ &\qquad\qquad\quad\cdot\Big{[}\sum_{k\in\mathbb{Z}}F^{*}_{+}({\nu}_{+,k})\wedge F^{*}_{+}(\bar{\nu}_{+,k})-F^{*}_{-}({\nu}_{-,k})\wedge F^{*}_{-}(\bar{\nu}_{-,k})\Big{]}\,d^{2}z\\ =&-\frac{\sqrt{-1}}{2}\int_{\Sigma}\,e^{\psi\circ F_{\pm}}\big{(}|(F_{\pm})_{z}|^{2}-|(F_{\pm})_{\bar{z}}|^{2}\big{)}\\ &\qquad\qquad\quad\cdot\Big{[}\sum_{k\in\mathbb{Z}}F^{*}_{+}({\nu}_{+,k})\wedge F^{*}_{+}(\bar{\nu}_{+,k})-F^{*}_{-}({\nu}_{-,k})\wedge F^{*}_{-}(\bar{\nu}_{-,k})\Big{]}\,d^{2}z\\ =&-\frac{\sqrt{-1}}{2}\Big{[}\mathrm{Mess}^{*}_{+}\int_{\mathbb{D}}e^{\psi}\,\Big{(}\sum_{k\in\mathbb{Z}}\nu_{+,k}\wedge\bar{\nu}_{+,k}\Big{)}\,d^{2}z\\ &\qquad\qquad\qquad\qquad\qquad-\mathrm{Mess}^{*}_{-}\int_{\mathbb{D}}e^{\psi}\Big{(}\sum_{k\in\mathbb{Z}}\nu_{-,k}\wedge\bar{\nu}_{-,k}\Big{)}\,d^{2}z\Big{]}.\end{split}$
Hence, the above equalities imply
@@
By Theorem 4.10, we can see that the symplectic form $\omega_{\mathrm{C}}$ does not depend on the choice of the basis $\{\nu_{\pm,k}:\,k\in\mathbb{Z}\}$.
@@
By the proof of Theorem 4.10, we can see that the pullback $\mathrm{Mess}_{\pm}^{*}(\omega_{\mathrm{WP}})$ has the following expression
(4.28) $\mathrm{Mess}^{*}_{\pm}(\omega_{\mathrm{WP}})=\frac{\sqrt{-1}}{2}\int_{\Sigma}(1-|\mu_{F}|^{2})\,e^{\phi}\,\Big{(}\sum_{k\in\mathbb{Z}}F^{*}_{\pm}(\nu_{\pm,k})\wedge F^{*}_{\pm}(\bar{\nu}_{\pm,k})\Big{)}\,d^{2}z.$
The closedness of this pullback 2-forms on $T^{*}T_{0}(1)$ just follows from the commutativity of $d$-differential and pullback map. But, it can be also proved as in the arguments in Lemma 2.7 of [24] or Theorem 7.4 of [15], whose original argument comes from [1]. For this argument, we need the fact that the Lie derivative of the density $(1-|\mu_{F}|^{2})e^{\phi}$ vanishes, which follows from Propositions 5.1 and 5.2.
@@
For the symplectic diffeomorphism $\mathrm{Mess}:T^{*}T_{0}(1)\to T_{0}(1)\times T_{0}(1)$,
@@
    $\mathrm{Mess}$ maps the Lagrangian subspace given by the zero section of $T^{*}T_{0}(1)$ to the diagonal subset in $T_{0}(1)\times T_{0}(1)$,
@@
    $\mathrm{Mess}$ maps the Lagrangian subspace given by the fiber $T^{*}_{0}T_{0}(1)$ to the anti-diagonal subset in $T_{0}(1)\times T_{0}(1)$.
@@
From the bottom part of the diagram (3.10), we have that $z_{\pm}=F_{\pm}\circ z$ in general. For the case of the Lagrangian subspace of the zero section of $T^{*}T_{0}(1)$, the map $F_{\pm}$ becomes trivial since the Hopf differential $\Phi$ is trivial. Hence, the disc $(\mathbb{D},z_{\pm})$ coincides with $(\mathbb{D},z)$, that is, $z_{\pm}=z$. In conclusion, $\mathrm{Mess}$ maps the Lagrangian of the zero section to the diagonal subset in $T_{0}(1)\times T_{0}(1)$. For the case of the Lagrangian subspace of the fiber $T_{0}^{*}T_{0}(1)$, the disc $(\mathbb{D},z)$ coincides with $(\mathbb{D},w)$, that is, $z_{\pm}=F_{\pm}$. Hence, the Beltrami differential of $z_{\pm}$ is the same as the Beltrami differential $\mu_{F_{\pm}}$. Moreover, $\mu_{F_{\pm}}=\pm\overline{\Phi}e^{-\phi}$. In conclusion, $\mathrm{Mess}$ maps the Lagrangian space of fiber $T_{0}^{*}T_{0}(1)$ to the anti-diagonal subset in $T_{0}(1)\times T_{0}(1)$. ∎

### v2

@@ 4.2. Symplectic forms
As described in Subsection 2.2, for a point $\mu\in T_{0}(1)$, the local coordinate system over an open neighborhood $V_{\mu}$ around $\mu$ is modeled on the Hilbert space $A_{2}(\mathbb{D}^{*})$. More precisely, fixing an orthonormal basis $\{\phi_{k}\}_{k\in\mathbb{Z}}$ of $A_{2}(\mathbb{D}^{*})$, the inverse images of $\phi_{k}$’s under the map given in (2.14) determine the coordinate system $\{\mu_{k}\}_{k\in\mathbb{Z}}$ over $V_{\mu}$. Thus, any point $\nu\in V_{\mu}$ can be expressed as
$\nu=\sum_{k\in\mathbb{Z}}\zeta_{k}\,\mu_{k},$
for small complex coefficients $\{\zeta_{k}\}_{k\in\mathbb{Z}}$.
Similarly, for a point $(\mu,\Phi)$ in the holomorphic cotangent bundle $T^{*}T_{0}(1)$, the local coordinate system on an open neighborhood around $(\mu,\Phi)$ is given by
$\{\zeta_{k},\eta_{k}\}_{k\in\mathbb{Z}}.$
Here the fiber coordinates $\{\eta_{k}\}_{k\in\mathbb{Z}}$ are also determined by the fixed orthonormal basis of $A_{2}(\mathbb{D}^{*})$, via the identification $T^{*}_{[\mu]}T_{0}(1)\cong A_{2}(\mathbb{D}^{*})$. Consequently, the differentials $(d\zeta_{k},d\eta_{k})_{k\in\mathbb{Z}}$ provide a basis for the cotangent space $T^{*}_{(\mu,\Phi)}(T^{*}T_{0}(1))$. Note that $d\zeta_{k}$ can be represented by harmonic Beltrami differentials, while $d\eta_{k}$ corresponds to holomorphic quadratic differentials.
We now examine how the quantities $d\zeta_{k}$ and $d\eta_{k}$ arise in the context of the deformations of the maximal conformal embeddings described in (4.1). To this end, let us begin by considering the following set up.
Let $F_{\pm}:\mathbb{D}_{z}\to\mathbb{D}_{\pm}$ be harmonic diffeomorphisms as given in the commutative diagram (4.14), and denote by $\mu_{\pm}$ the Beltrami differentials representing $\mathbb{D}_{\pm}$ respectively. As described in Subsection 2.2, we fix an orthonormal basis in $H^{-1,1}(\mathbb{D})$ and obtain the corresponding orthonormal basis for $T_{[\mu]}T_{0}(1)$ by applying right translations: $D_{0}R_{[\mu]}(H^{-1,1}(\mathbb{D}))$ for any $[\mu]\in T_{0}(1)$. We denote by $\{\nu_{\pm,k}\}_{k\in\mathbb{Z}}$ the resulting orthonormal basis of the tangent space $T_{[\mu_{\pm}]}T_{0}(1)$.
Given a harmonic Beltrami differential $\nu_{\pm,k}$ on $\mathbb{D}_{\pm}$, we solve the Beltrami equation
(4.19) $(h^{\epsilon}_{\pm})_{\bar{z}}=\epsilon\nu_{\pm,k}(h^{\epsilon}_{\pm})_{z}\qquad\text{where}\quad z=z_{\pm},$
to obtain a quasi-conformal map $h^{\epsilon}_{\pm}:\mathbb{D}_{\pm}\to\mathbb{D}_{\pm}^{\epsilon}$. Note that $h^{\epsilon}_{\pm}$ and $\mathbb{D}^{\epsilon}_{\pm}$ depend on index $k$, though we supress this dependence in the notation for simplicity.
Given the two deformed discs $\mathbb{D}^{\epsilon}_{\pm}$ depending on the same index $k$, we construct a maximal conformal embedding $\sigma^{\epsilon}:\mathbb{D}^{\epsilon}\to\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}$, as the inverse of the $\mathrm{Mess}$ map, such that the resulting induced Gauss maps coincide with the given harmonic maps
$F_{\pm}^{\epsilon}:\mathbb{D}^{\epsilon}\to\mathbb{D}^{\epsilon}_{\pm}.$
Correspondingly, there exists a unique quasi-conformal map
$f^{\epsilon}:\mathbb{D}\to\mathbb{D}^{\epsilon}$
satisfying the relations
@@
This follows from the fact that the deformed disk $\mathbb{D}^{\epsilon}$ is realized as the graph of the map $F^{\epsilon}_{+}\circ(F^{\epsilon}_{-})^{-1}$ in $\mathbb{D}_{+}^{\epsilon}\times\mathbb{D}_{-}^{\epsilon}$, and each point in $\mathbb{D}^{\epsilon}$ is determined by its coordinates via the maps $F^{\epsilon}_{\pm}$. Note that the both the map $f^{\epsilon}$ and the deformed disc $\mathbb{D}^{\epsilon}$ depend on the index $k$, although we omit this dependence in the notation for simplicity.
We now define $\nu^{\epsilon}_{k}$ as the Beltrami differential on $\mathbb{D}$ corresponding to $f^{\epsilon}_{k}$, that is,
$(f^{\epsilon}_{k})_{\bar{z}}=\nu^{\epsilon}_{k}\,(f^{\epsilon}_{k})_{z}.$
Let $\nu_{k}$ denotes the harmonic Beltrami differential given by the derivative of $\nu_{k}^{\epsilon}$ at $\epsilon=0$. By construction, the basis $\{\nu_{k}\}_{k\in\mathbb{Z}}$ coincides with the one determined by $\{\nu_{\pm,k}\}$ via the right translations. We define the variation of $\mu$ along $\nu_{k}$ by
@@
where the second equality follows from equation(4.16). By (4.7), we observe that $F^{*}_{\pm}(\nu_{\pm,k})$ is determined by $\nu_{k}$ and $\mu_{F}=\mu_{F_{+}}$. Similarly, the variation of $\Phi$ along $\nu_{k}$ is defined by
@@
where the second equality follows from equation(4.15).
@@
The differential of the $\mathrm{Mess}$ map
$\mathrm{Mess}:T^{*}T_{0}(1)\to T_{0}(1)\times T_{0}(1)$
is an isomorphism.
@@
By the proof of Theorem 3.10, the $\mathrm{Mess}$ map is bijective, implying that its inverse $\mathrm{Mess}^{-1}$ is well-defined. We now show that the differential of $\mathrm{Mess}^{-1}$ is an isomorphism. First, we observe that the differential of $\mathrm{Mess}^{-1}$ is given by (4.21) and (4.22). To analyze these maps, we decompose the tangent space $T(T_{0}(1)\times T_{0}(1))$ at $z_{\pm}|_{\mathbb{S}^{1}}$ into the tangent space of the diagonal and the tangent space of the anti-diagonal. Along the tangent space of the diagonal, we have $F^{*}_{+}(\nu_{+,k})=F^{*}_{-}(\nu_{-,k})$. Thus, the component of the differential of $\mathrm{Mess}^{-1}$ given in (4.22) vanishes over this subspace. If the component of the differential of $\mathrm{Mess}^{-1}$ given in (4.21) has a nontrivial kernel, then there must exist a harmonic Beltrami differential $\nu_{+}$ such that
$F^{*}_{+}(\nu_{+})+F^{*}_{+}(\bar{\nu}_{+})\mu_{F}^{2}=0.$
Combining this with its complex conjugate equation, we obtain $F^{*}_{+}(\nu_{+})(1-|\mu_{F}|^{4})=0$, which can not occur since $|\mu_{F}|<1$ and $F_{+}$ is a diffeomorphism. Hence, the component of the differential of $\mathrm{Mess}^{-1}$ given in (4.21) is injective.
To show that this component is surjective over the tangent space of the diagonal, let $\nu$ be a harmonic Beltrami differential representing a tangent vector $\delta\mu$, Then, there exists a Beltrami differential $\nu_{+}$ satisfying
$\nu-\bar{\nu}\mu_{F}^{2}=(1-|\mu_{F}|^{2})F_{+}^{*}(\nu_{+}),$
by the fact $F$ is a diffeomorphism. This implies
@@
Since $\nu$ can be written in terms of the basis $\nu_{k}$’s, the same holds for $\nu_{+}$ in terms of the basis $\nu_{+,k}$’s. Thus, this component is surjective over the tangent space of the diagonal.
Similarly, we can show that the differential of $\mathrm{Mess}^{-1}$ given in (4.22) is bijective over the tangent space of the anti-diagonal. Therefore, combining these results, we conclude that the differential of $\mathrm{Mess}:T^{*}T_{0}(1)\to T_{0}(1)\times T_{0}(1)$ is an isomorphism. This completes the proof. ∎
By Proposition 4.7, we conclude that the families $\{\delta_{k}\mu\}_{k\in\mathbb{Z}}$ and $\{\delta_{k}\Phi\}_{k\in\mathbb{Z}}$ together form a basis for the cotangent bundle of $T^{*}T_{0}(1)$ at $(\mu,\Phi)$. Therefore, they represent the realization of the basis $\{d\zeta_{k},d\eta_{k}\}_{k\in\mathbb{Z}}$ in the context of a deformation of the maximal conformal embedding (4.1).
Remark 4.8.
Strictly speaking, the basis $\{d\zeta\}_{k\in\mathbb{Z}}$ is realized by the duals of the variations $\{\delta_{k}\mu=\nu_{k}\}_{k\in\mathbb{Z}}$ rather than the variations themselves. However, for simplicity of notation, we will not distinguish between these in the subsequent constructions.
Based on this identification, the complex canonical symplectic form $\omega_{\mathbb{C}}$ on $T^{*}T_{0}(1)$ admits the following expression:
$\omega_{\mathbb{C}}=\int_{\mathbb{D}}\ \sum_{k\in\mathbb{Z}}\delta_{k}\mu\wedge\delta_{k}\Phi\ d^{2}z.$
Moreover, its imaginary part, referred to as the canonical symplectic form, is given by
(4.23) $\omega_{\mathrm{C}}=\int_{\mathbb{D}}\Big{(}\sum_{k\in\mathbb{Z}}\sqrt{-1}\delta_{k}\Phi\wedge\delta_{k}\mu-\sqrt{-1}\delta_{k}\overline{\Phi}\wedge\delta_{k}\bar{\mu}\Big{)}\,d^{2}z.$
Note that the wedge product $\wedge$ is applied to the vector-valued expressions $F^{*}_{\pm}(\nu_{\pm,k})$ appearing on the right hand sides of (4.21) and (4.22). The construction of $\omega_{\mathrm{C}}$ is globally well-defined, since all the local expressions are compatible through right translations from a neighborhood $V_{0}$ of the origin in $T_{0}(1)$. While the definition of $\omega_{\mathrm{C}}$ may seem to depend on the choice of basis $\{\nu_{\pm,k}|\,k\in\mathbb{Z}\}$, we will later show that $\omega_{\mathrm{C}}$ is in fact independent of this choice (see Remark 4.11).
@@
The following equality holds:
(4.24) $\begin{split}\omega_{\mathrm{C}}=&-\frac{\sqrt{-1}}{2}\int_{\mathbb{D}}(1-|\mu_{F_{+}}|^{2})\,e^{\phi}\\ &\qquad\qquad\cdot\,\sum_{k\in\mathbb{Z}}\Big{(}F^{*}_{+}(\nu_{+,k})\wedge F^{*}_{+}(\bar{\nu}_{+,k})-F^{*}_{-}(\nu_{-,k})\wedge F^{*}_{-}(\bar{\nu}_{-,k})\Big{)}\,d^{2}z.\end{split}$
@@
$\qquad\qquad\qquad\qquad\qquad\qquad\quad\wedge\Big{(}(F^{*}_{+}(\bar{\nu}_{+,k})+F^{*}_{-}(\bar{\nu}_{-,k}))+(F^{*}_{+}({\nu}_{+,k})+F^{*}_{-}({\nu}_{-,k})\bar{\mu}_{F}^{2}\Big{)}.$
Rearranging the terms, we obtain
@@
Note that for the orthonormal basis $\{\nu_{\pm,k}\}_{k\in\mathbb{Z}}$ with respect to the Weil-Petersson inner product on $T_{[\mu_{\pm}]}T_{0}(1)$, the Weil-Petersson symplectic form $\omega_{\mathrm{WP}}$ on $T_{0}(1)$ is given by
@@
The map
$\mathrm{Mess}:T^{*}T_{0}(1)\to T_{0}(1)\times T_{0}(1)$
is a symplectic diffeomorphism such that
@@
Here, $\mathrm{Mess}_{\pm}:=\pi_{\pm}\circ\mathrm{Mess}$, where $\pi_{\pm}$ denotes the projection map from $T_{0}(1)\times T_{0}(1)$ onto the first or second factor, respectively.
@@
By Proposition 4.9, we have
(4.27) $\begin{split}\omega_{\mathrm{C}}=&-\frac{\sqrt{-1}}{2}\int_{\mathbb{D}}(1-|\mu_{F}|^{2})\,e^{\phi}\\ &\qquad\qquad\quad\cdot\Big{[}\sum_{k\in\mathbb{Z}}F^{*}_{+}({\nu}_{+,k})\wedge F^{*}_{+}(\bar{\nu}_{+,k})-F^{*}_{-}({\nu}_{-,k})\wedge F^{*}_{-}(\bar{\nu}_{-,k})\Big{]}\,d^{2}z\\ =&-\frac{\sqrt{-1}}{2}\int_{\mathbb{D}}\,e^{\psi\circ F_{\pm}}\big{(}|(F_{\pm})_{z}|^{2}-|(F_{\pm})_{\bar{z}}|^{2}\big{)}\\ &\qquad\qquad\quad\cdot\Big{[}\sum_{k\in\mathbb{Z}}F^{*}_{+}({\nu}_{+,k})\wedge F^{*}_{+}(\bar{\nu}_{+,k})-F^{*}_{-}({\nu}_{-,k})\wedge F^{*}_{-}(\bar{\nu}_{-,k})\Big{]}\,d^{2}z\\ =&-\frac{\sqrt{-1}}{2}\Big{[}\mathrm{Mess}^{*}_{+}\int_{\mathbb{D}}e^{\psi}\,\Big{(}\sum_{k\in\mathbb{Z}}\nu_{+,k}\wedge\bar{\nu}_{+,k}\Big{)}\,d^{2}z\\ &\qquad\qquad\qquad\qquad\qquad-\mathrm{Mess}^{*}_{-}\int_{\mathbb{D}}e^{\psi}\Big{(}\sum_{k\in\mathbb{Z}}\nu_{-,k}\wedge\bar{\nu}_{-,k}\Big{)}\,d^{2}z\Big{]}.\end{split}$
Thus, we obtain
@@
By Theorem 4.10, we can see that the canonical symplectic form $\omega_{\mathrm{C}}$ does not depend on the choice of the basis $\{\nu_{\pm,k}:\,k\in\mathbb{Z}\}$.
@@
For any harmonic Beltrami differentials $\mu_{1},\mu_{2}$ representing vectors in the tangent space of $T_{\nu}T_{0}(1)\cong H^{-1,1}(\mathbb{D})$ at $[\nu]\in T_{0}(1)$, we easily obtain the inequality
$\big{|}\,\omega_{\mathrm{WP}}(\mu_{1},\mu_{2})\,\big{|}^{2}\leq||\mu_{1}||_{2}\cdot||\mu_{2}||_{2}<\infty$
where $||\cdot||_{2}$ denotes the $L^{2}$-norm on $H^{-1,1}(\mathbb{D})$. This observation, together with Theorem 4.10, implies that the canonical symplectic form $\omega_{\mathrm{C}}$ on $T^{*}T_{0}(1)$ satisfies the same boundedness property.
Remark 4.13.
From the proof of Theorem 4.10, we observe that the pullback $\mathrm{Mess}_{\pm}^{*}(\omega_{\mathrm{WP}})$ has the following expression:
(4.28) $\mathrm{Mess}^{*}_{\pm}(\omega_{\mathrm{WP}})=\frac{\sqrt{-1}}{2}\int_{\mathbb{D}}(1-|\mu_{F}|^{2})\,e^{\phi}\,\Big{(}\sum_{k\in\mathbb{Z}}F^{*}_{\pm}(\nu_{\pm,k})\wedge F^{*}_{\pm}(\bar{\nu}_{\pm,k})\Big{)}\,d^{2}z.$
The closedness of these pullback 2-forms on $T^{*}T_{0}(1)$ follows from the commutativity of the exterior differential $d$ with the pullback operation. Alternatively, it can be established using similar arguments as in Lemma 2.7 of [27] or Theorem 7.4 of [18], which in turn originate from [1]. A key component of this argument is the fact that the Lie derivative of the density $(1-|\mu_{F}|^{2})e^{\phi}$ vanishes, a result that follows from Propositions 5.1 and 5.2.
@@
For the symplectic diffeomorphism $\mathrm{Mess}:T^{*}T_{0}(1)\to T_{0}(1)\times T_{0}(1)$, the following properties hold:
@@
    The $\mathrm{Mess}$ map sends the Lagrangian subspace given by the zero section of $T^{*}T_{0}(1)$ to the diagonal subset in $T_{0}(1)\times T_{0}(1)$,
@@
    The $\mathrm{Mess}$ map sends the Lagrangian subspace given by the fiber $T^{*}_{0}T_{0}(1)$ to the anti-diagonal subset in $T_{0}(1)\times T_{0}(1)$.
@@
The first claim follows directly from Remark 3.12. For the second claim concerning the Lagrangian subspace of the fiber $T_{0}^{*}T_{0}(1)$, we observe that the disc $\mathbb{D}_{z}$ coincides with $\mathbb{D}_{w}$, meaning that $z_{\pm}=F_{\pm}$. Consequently, the Beltrami differentials associated with $z_{\pm}$ are identical to the Beltrami differential $\mu_{F_{\pm}}$. Furthermore, we have $\mu_{F_{\pm}}=\pm\overline{\Phi}e^{-\phi}$. Thus, $\mathrm{Mess}$ maps the Lagrangian subspace corresponding to the fiber of $T_{0}^{*}T_{0}(1)$ to the anti-diagonal subset in $T_{0}(1)\times T_{0}(1)$. ∎